Imagine that you are in an argument with your sibling or friend about the size of the slice of pizza that you were both served. They say that it was unfair because you got a bigger piece. What information would you need to know about each slice of pizza in order to calculate which one had a larger area? Explain. (hint: remember the formula for the area of a sector of a circle!! )


1
Expert's answer
2021-05-10T18:33:12-0400

Let r be the radius of the pizza.

The angled subtends at the center.


So That we can calculate the area as-


"\text{ Area }=\dfrac{\pi r^2 \theta}{360}"


where, "\theta=" Angle subtends by pizza arc